OpenStudy (anonymous):

resolve into partial fraction:((2x)/((x^2-1)(x-1)))

OpenStudy (anonymous):

\[\frac{2 x}{(x-1) \left(x^2-1\right)}=-\frac{1}{2 (x+1)}+\frac{1}{2 (x-1)}+\frac{1}{(x-1)^2} \]